### v1

Abstract
Large language model (LLM) queries are predominantly processed by frontier models in centralized cloud infrastructure. Rapidly growing demand strains this paradigm, and cloud providers struggle to scale infrastructure at pace. Two advances create an opportunity to rethink this paradigm: small, local LMs ($\leq$ 20B active parameters) now achieve competitive performance to frontier models on many tasks, and local accelerators (e.g., Apple M4 Max) can host these models at interactive latencies. This raises the question: can local inference viably redistribute demand from centralized infrastructure? Answering this requires measuring both whether local LMs can accurately answer real-world queries and whether they can do so efficiently enough to be practical on power-constrained devices (i.e., laptops). We propose intelligence per watt (IPW), task accuracy divided by unit of power, as a unified metric for assessing both the capability and efficiency of local inference across model-accelerator configurations. We conduct a large-scale empirical study across 20+ state-of-the-art local LMs, 8 hardware accelerators (local and cloud), and a representative subset of LLM traffic: 1M real-world single-turn chat and reasoning queries. For each query, we measure accuracy (local LM win rate against frontier models), energy consumption, latency, and power. Our analysis reveals three key findings. First, that local LMs can successfully answer $88.7\%$ of single-turn chat and reasoning queries with accuracy varying by domain. Second, longitudinal analysis from 2023-2025 shows progress in local inference viability: IPW improved $5.3\times$, driven by both algorithmic advances and accelerator improvements, with locally-serviceable query coverage increasing from 23.2% to 71.3%. Third, local accelerators achieve at least 1.4× lower IPW than cloud accelerators running identical models, revealing significant headroom for local accelerator optimization These findings demonstrate that local inference can meaningfully redistribute demand from centralized infrastructure for a substantial subset of queries, with IPW serving as the critical metric for tracking this transition. We release our IPW profiling harness to enable intelligence-per-watt benchmarking as local LMs and accelerators evolve.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
Large language model (LLM) queries are predominantly processed by frontier models deployed in centralized cloud infrastructure (OpenAI, 2025; Alvarez & Marsal, 2025). This centralized approach faces mounting resource constraints as inference workloads scale from billions to trillions of queries daily (Alvarez & Marsal, 2025). History suggests an alternative path forward. From 1946-2009, computing efficiency (performance-per-watt) doubled every 1.5 years (Koomey et al., 2010), enabling a redistribution of computing workloads from data center mainframes to personal computers. This transition occurred when efficiency improvements enabled computing to meet user needs within personal device power constraints, not when PCs surpassed mainframes in raw performance.
Two converging trends suggest a similar inflection point may be emerging for LLM inference. First, recent advances have produced local LMs: small models ($\leq 20B$ active parameters) such as Qwen3 (Team, 2025), Llama3.1 (Grattafiori et al., 2024), and gpt-oss (Agarwal et al., 2025) that achieve competitive performance on many benchmarks while requiring less energy and compute than larger, frontier models (Agarwal et al., 2025). Second, local accelerators (e.g., Apple M4 Max, AMD Ryzen AI) now have sufficient memory capacity and compute throughput to host these models with interactive latencies (Apple, 2024). This raises the question: Can local inference viably redistribute demand from centralized infrastructure?
Answering this requires measuring two factors: the capability of local LMs to accurately respond to a subset of real-world queries, and the efficiency with which local accelerators convert power into useful computation. To asses this, we need a unified metric that captures both the intelligence delivered (model capability) and the energy required (accelerator efficiency). We introduce intelligence per watt (IPW): task accuracy per unit of power consumption. IPW directly measures the fundamental tradeoff facing local inference: achieving sufficient task performance within constrained power budgets. This metric enables systematic comparison across model-accelerator configurations and quantifies efficiency gains from model architecture innovations (Team, 2025; OpenAI, 2025; Team et al., 2025a; IBM Research, 2025), post-training techniques (Hinton et al., 2015; Ouyang et al., 2022; Shao et al., 2024a; Dettmers et al., 2022), and accelerator improvements (Median-Group, 2019; NVIDIA Corporation, 2025b; AMD, 2025).
To evaluate the viability of local inference and measure progress in IPW, we conduct a large-scale empirical study addressing three questions:
  • Q1:
    What fraction of current inference queries can be solved by local LMs on local accelerators, and how has this changed over time?
  • Q2:
    How has intelligence per watt improved across successive generations of local models and accelerators, and what are the relative contributions of model versus accelerator advances?
  • Q3:
    What resource savings (e.g. compute, energy, dollar cost) are possible by distributing workloads across local and cloud infrastructure?
Our study evaluates 20+ local LMs across 8 hardware accelerators on 1M queries spanning naturalistic user conversations (Deng et al., 2024), general reasoning tasks (Yuan et al., 2025), and standardized benchmarks measuring knowledge breadth (MMLU Pro (Wang et al., 2024b)) and expert-level reasoning (SuperGPQA (Team et al., 2025b)). We focus on single-turn interactions because they constitute a substantial portion of LLM usage (Deng et al., 2023; Wang et al., 2024a; Shirey, 2025). We compare state-of-the-art local LMs from October 2025—Qwen3 (Team, 2025), gpt-oss (Agarwal et al., 2025), Gemma3 (Team et al., 2025a), and IBM Granite4 (IBM Research, 2025)—alongside 2023-2024 models (Mixtral-8x7B, Llama-3.1-8B) on NVIDIA, AMD, and Apple accelerators. For each query, we measure accuracy, latency, energy, compute, cost, and memory, enabling systematic analysis of how intelligence per watt has evolved from 2023-2025 (Section 4). We release our hardware-agnostic profiling harness to support reproducible efficiency benchmarking as new models and accelerators emerge.
Figure 1: Intelligence per Watt: A Study of Local Intelligence Efficiency. We present the first systematic study of local AI inference efficiency across models, hardware, and real-world workloads. (Left) Intelligence efficiency is defined as task accuracy per unit of power, capturing both capabilities delivered and energy consumed. (Left-Middle) We conduct comprehensive performance profiling across 20+ state-of-the-art local LMs ($\leq 20B$ active parameters), diverse hardware accelerators (Apple, NVIDIA, AMD), multiple performance metrics, and 1M+ real-world queries spanning chat and reasoning tasks. (Right-Middle) Local LM capabilities are improving rapidly: win/tie rate versus frontier models increases from $23.2\%$ (2023) to $71.3\%$ (2025)—a $3.1\times$ improvement in accuracy, demonstrating that local models can accurately handle significant portions of single-turn chat and reasoning queries. (Right) Intelligence per watt improves $5.3\times$ from 2023–2025, driven by advances in both model architectures and hardware accelerators, with local accelerators showing $1.5\times$ efficiency headroom compared to enterprise-grade systems.
To address Q1, we study local LM coverage: the fraction of queries correctly answered (as measured by win-rate over a frontier LM) by at least one local LM in our study. Our analysis reveals that $88.7\%$ of queries can be successfully handled by small local models as of October 2025, with coverage varying by domain—exceeding $90\%$ for creative tasks (e.g., Arts & Media) but dropping to $68\%$ for technical fields (e.g., Architecture & Engineering) (Figure 7). Longitudinal analysis shows consistent improvement: the best local LM matched frontier model quality on $23.2\%$ of queries in 2023, $48.7\%$ in 2024, and $71.3\%$ in 2025—a $3.1\times$ increase over two years (Table 2).
To address Q2, we track IPW improvements across successive generations of models and accelerators from 2023-2025, decomposing gains by isolating model and accelerator contributions (Table 2). Holding accelerator fixed at NVIDIA H100, model improvements from Mixtral-8x7B (2023) to gpt-oss-120b (2025) yield a $3.1\times$ gain in accuracy per watt; conversely, holding the model fixed at gpt-oss-120b, accelerator improvements from H100 to Blackwell deliver a $1.7\times$ gain. Combined, these compounding improvements produce a $5.3\times$ overall increase (Figure 5).
To address Q3, we quantify the potential resource savings from treating cloud and local infrastructure as complementary resources and routing queries to local LMs when capable, otherwise to frontier models. Oracle routing (perfect assignment of each query to the smallest capable model) could reduce energy consumption by $80.4\%$, compute by $77.3\%$, and cost by $73.8\%$ versus cloud-only deployment (Figure 6), representing the theoretical maximum efficiency gains. Notably, routing systems need not achieve perfect accuracy to realize substantial savings while maintaining task quality. A routing system with $80\%$ accuracy (correctly assigning $80\%$ of queries to local vs. cloud) captures $80\%$ of theoretical maximum gains, achieving $64.3\%$ energy reduction, $61.8\%$ compute reduction, and $59.0\%$ cost reduction with no degradation in answer quality.
Our work makes three primary contributions. (1) We introduce intelligence per watt as a unified metric for evaluating local inference viability, and conduct the first large-scale empirical study measuring its evolution across 1M+ queries, 20+ models, and 8 hardware accelerators spanning 2023-2025. (2) We demonstrate that local LMs can accurately answer $88.7\%$ of single-turn chat and reasoning queries, with intelligence per watt improving $5.3\times$ over two years through compounding model ($3.2\times$) and hardware ($1.7\times$) advances. (3) We show that hybrid local-cloud systems achieve $40-65\%$ reductions in energy, compute, and cost with realistic routing accuracy while maintaining answer quality. Together, these findings establish local inference as a practical complement to centralized infrastructure whose viability continues expanding. We release our IPW profiling harness to facilitate systematic intelligence-per-watt benchmarking as local LMs and accelerators evolve.
2 Related Work
Our work is inspired by prior work on local AI and local-cloud inference system studies. See App. A for an extended related work.
Inference Workload Demand. The demand for AI inference is growing exponentially, placing unprecedented strain on computational infrastructure. Recent industry analyses project that global data center capacity could nearly triple by 2030, reaching 156-219 GW, with approximately 70% of new demand driven by AI workloads (McKinsey & Company, 2025). Meeting this demand will require \$5.2-\$7.9 trillion in capital expenditures across computing hardware, power infrastructure, and data center construction, depending on whether growth follows constrained, continued, or accelerated scenarios (McKinsey & Company, 2025). In the U.S. alone, AI systems could require over 50 GW of power capacity by 2030, approaching 5% of total U.S. power generation capacity (You et al., 2025), while global projections suggest AI data centers may need 68-100 GW by 2027-2030 (Pilz et al., 2025a; b).
[…]
Local-Cloud Inference Systems Recent work explores collaborative protocols for splitting generation between local and cloud LMs. Prior approaches focus on token- and layer-level collaboration: Minions proposes a communication protocol where an on-device LM handles lightweight processing while a cloud LM performs high-level reasoning (Narayan et al., 2025), speculative decoding employs draft model verification (Miao et al., 2023; Xu et al., 2025), and systems like SLED, HAT, and CE-CoLLM introduce edge-cloud partitioning with early-exit mechanisms (Li et al., 2025; Xie et al., 2025; Jin & Wu, 2025). These techniques have been extended beyond language models to diffusion models (Yan et al., 2024), with collaboration mechanisms broadly categorized as pipeline, routing, auxiliary, distillation, and fusion strategies (Chen et al., 2025). Complementary infrastructure work addresses efficient GPU resource management for concurrent model serving (Xiang et al., 2025). Recent benchmarks have expanded beyond academic tasks to measure AI performance on real-world economically valuable work across professional domains (OpenAI, 2025a; Mercor, 2025; Snorkel AI, 2025). Evaluating local-cloud systems on such benchmarks is critical: if these systems can maintain quality on economically important tasks while offloading work from cloud to local devices, the infrastructure savings compound on the workloads that drive actual productivity and economic value. In contrast to this line of work, we seek to better understand the current limitations and improvement trajectory of local inference as a means of efficiently redistributing LLM traffic between local and cloud resources.
3 Preliminaries
We formalize local and cloud inference infrastructure and introduce metrics for measuring intelligence efficiency.
Inference Infrastructure: Queries, Models and Accelerators. We consider an inference infrastructure serving a stream of user queries $\mathcal{Q}=\{q_{1},q_{2},\ldots,q_{n}\}$, where each query $q_{i}$ represents a user-generated request (e.g., chat messages, reasoning tasks). Let $\mathcal{M}_{\text{local}}=\{m_{1},\ldots,m_{k}\}$ denote a set of local LMs with $\leq 20B$ active parameters each, and $\mathcal{M}_{\text{cloud}}=\{M_{1},\ldots,M_{\ell}\}$ denote frontier LMs with $\geq 100B$ parameters. Similarly, let $\mathcal{H}_{\text{local}}$ represent local accelerators (e.g., Apple M4, AMD Ryzen) and $\mathcal{H}_{\text{cloud}}$ represent cloud accelerators (e.g., NVIDIA H200, AMD MI300X).
Inference Serving: Local and Cloud. We distinguish between two inference paradigms: local inference, where queries are processed by models $m\in\mathcal{M}_{\text{local}}$ on accelerator $h\in\mathcal{H}_{\text{local}}$, and cloud inference, where queries are processed by models $M\in\mathcal{M}_{\text{cloud}}$ on accelerator $H\in\mathcal{H}_{\text{cloud}}$. A routing function $r:\mathcal{Q}\rightarrow\mathcal{M}_{\text{local}}\cup\mathcal{M}_{\text{cloud}}$ determines the assignment of each query to either a local or cloud model. In this work we consider a setup where queries can be routed between multiple local models $m\in\mathcal{M}_{\text{local}}$ hosted on a local accelerator $h\in\mathcal{H}_{\text{local}}$ with support for small language models (up to $20B$ active parameters), and cloud models $M\in\mathcal{M}_{\text{cloud}}$ hosted on a cloud accelerator $H\in\mathcal{H}_{\text{cloud}}$ with support for large-scale frontier models (at least 100B parameters).
Intelligence Efficiency Metrics. We introduce a family of metrics to quantify how efficiently inference systems convert energy into useful computation. For a model-accelerator pair $(m,h)$, let $\text{acc}(m,q)$ denote the accuracy of model $m$ on query $q$, $\text{ppl}(m,q)$ denote the perplexity, $P(m,h,q)$ denote the average power consumption (in watts) during inference for query $q$, and $\tau(m,h,q)$ denote the total latency (in seconds) for generating the response, including both prefill and decoding phases.
We define four complementary efficiency metrics:
Power-based metrics measure efficiency relative to instantaneous power draw:
  • Accuracy per watt: $\text{APW}(m,h)=\frac{\mathbb{E}_{q\sim\mathcal{Q}}[\text{acc}(m,q)]}{\mathbb{E}_{q\sim\mathcal{Q}}[P(m,h,q)]}$
  • Perplexity per watt:
    $\text{PPW}(m,h)=\frac{1}{\mathbb{E}_{q\sim\mathcal{Q}}[\text{ppl}(m,q)]\cdot\mathbb{E}_{q\sim\mathcal{Q}}[P(m,h,q)]}$
Energy-based metrics measure efficiency relative to total energy consumed per query:
  • Accuracy per joule:
    $\text{APJ}(m,h)=\frac{\mathbb{E}_{q\sim\mathcal{Q}}[\text{acc}(m,q)]}{\mathbb{E}_{q\sim\mathcal{Q}}[P(m,h,q)\cdot\tau(m,h,q)]}$
  • Perplexity per joule:
    $\text{PPJ}(m,h)=\frac{1}{\mathbb{E}_{q\sim\mathcal{Q}}[\text{ppl}(m,q)]\cdot\mathbb{E}_{q\sim\mathcal{Q}}[P(m,h,q)\cdot\tau(m,h,q)]}$
where $P(m,h,q)\cdot\tau(m,h,q)$ represents the energy consumption (in joules) for processing query $q$.
Power-based metrics (APW, PPW) capture the instantaneous efficiency of the inference system, reflecting the hardware’s ability to deliver performance at a given power draw. Energy-based metrics (APJ, PPJ) capture the total efficiency per query, accounting for both power consumption and generation latency. Together, these metrics provide a comprehensive view of inference efficiency: intelligence per watt quantifies the steady-state efficiency of model-accelerator pairs, while intelligence per joule quantifies the end-to-end efficiency from a user’s perspective, including the time cost of generation.
4 Dataset and Profiling Harness
In this section, we provide details on the dataset selection and profiling harness.
4.1 Dataset Selection
Query Curation We curate over $1M$ queries across four complementary benchmarks designed to measure both naturalistic deployment scenarios and controlled capability assessment. To ensure our findings about local inference efficiency generalize across task distributions, we combine naturalistic queries that reflect real-world LLM usage patterns with standardized benchmarks that enable systematic evaluation of knowledge breadth and reasoning capabilities across diverse domains.
For naturalistic chat tasks, we source queries from Wildchat (Deng et al., 2024): a dataset of $1M$ real ChatGPT prompts, spanning 1 month of user traffic. For general reasoning tasks, we source queries from NaturalReasoning (Yuan et al., 2025), which provides approximately 1.2 million reasoning-focused queries spanning diverse domains including mathematics, physics, and chemistry. For standardized knowledge evaluation, we use MMLU Pro (Wang et al., 2024b): an enhanced version of MMLU with increased difficulty (10 vs. 4 answer choices) and improved robustness to prompt variations, measuring multi-domain knowledge understanding. For expert-level reasoning across specialized disciplines, we evaluate on SuperGPQA (Team et al., 2025b): a comprehensive benchmark spanning 285 graduate-level disciplines with emphasis on technical domains and specialized fields underrepresented in typical evaluations (e.g., light industry, agriculture, service sciences).
We perform robust data cleaning and filtering (see App. B.1 for details) on each dataset before sampling queries: $500K$ from Wildchat, $500K$ from NaturalReasoning, $12K$ from MMLU Pro, and $26.5K$ from SuperGPQA (see Table 1). To provide finer-grained labels beyond task categories, we use gpt-4o-mini to annotate each query with a category from the Anthropic Economic Index (Handa et al., 2025), which maps AI queries to occupations in the U.S. Department of Labor’s O*NET. We consider 22 categories in total, spanning “Architecture and Engineering” to “Healthcare Support”. The full list appears in App. B.1 (Table 5) along with a breakdown of category representation in the dataset.
Dataset Origin Category $|N|$
Wildchat Chat Queries 500K
NaturalReasoning Reasoning Queries 500K
MMLU Pro Knowledge Evaluation 12K
SuperGPQA Graduate-Level Reasoning 26.5K
Category Items
Model Families Qwen3, gpt-oss, Gemma, IBM Granite 4.0
Accelerators NVIDIA A100, NVIDIA H200
NVIDIA GH200, NVIDIA B200
NVIDIA Quadro RTX 6000, NVIDIA RTX 6000 Ada
AMD MI300X, Apple M4 Max, Sambanova SN40L
Table 1: Dataset Overview. (Top) Query composition with split sizes. (Bottom) Models and Hardware Accelerators.
Hardware Accelerators We run inference on six different AI accelerator systems: the NVIDIA A100 40 GB SXM4 (Ampere) (NVIDIA, 2021), NVIDIA H200 SXM (Hopper) (NVIDIA, 2024), NVIDIA GH200 Grace Hopper Superchip (NVIDIA Corporation, 2024), NVIDIA B200 (Blackwell) (NVIDIA Corporation, 2025a), NVIDIA Quadro RTX 6000 (NVIDIA Corporation, 2019), NVIDIA RTX 6000 Ada (NVIDIA Corporation, 2023), AMD Instinct MI300X (CDNA 3, OAM) (AMD, 2023), Sambanova SN40L (SambaNova Systems, 2022) and Apple Mac Studio (M4 Max) (Apple, 2024). These systems were chosen because of their different memory capacities (ranging from 40 GB to 768 GB), memory bandwidth (from 546 GB/s to 8 TB/s), and power consumption (145W to 1000W) (see Table 9 for more details).
Models We collect model generations over the Qwen3 (Team, 2025), gpt-oss (OpenAI, 2025), Gemma3 (Team et al., 2025a), and IBM Granite 4.0 (IBM Research, 2025) model families. For the Qwen3 family, we use Qwen3-4B, Qwen3-8B, Qwen3-14B, Qwen3-32B, and Qwen3-235B. For GPT-OSS, we consider the gpt-oss-20b and gpt-oss-120b models. For the Gemma3 family, we use Gemma3 1B Instruct, Gemma3 4B Instruct, and Gemma3 12B Instruct models. For the IBM Granite 4.0 family, we use granite-4.0-h-micro, granite-4.0-h-tiny, granite-4.0-h-small, and granite-4.0-h-tiny models. We also benchmark against state-of-the-art cloud models as of October 2025, including Claude Sonnet 4.5 (Anthropic, 2025), Gemini 2.5 Pro (Comanici et al., 2025), and GPT-5 (2025-08-07) (OpenAI, 2025b). For our longitudinal analysis, we evaluate Mixtral-8x7B (Jiang et al., 2024) as well as Llama3.1-8B (Jiang et al., 2024). For each model, we generate responses across all dataset queries on each of the hardware backends. Full details of inference hyperparameters can be found in App. B.1.
Metrics For each (query, model, hardware) triple, we collect labels for accuracy, as well as a wide range of efficiency measurements: latency, throughput, time-to-first-token (TTFT), energy consumption, and more (see Table 8 for a full list of metrics). We use LLM-as-a-judge (see App. B.1 for the respective prompts) to score generated responses against reference answers. For Wildchat, reference answers are responses from Qwen3-235B, the SOTA open-source model on LMArena (as of August 2025) (Chiang et al., 2024). For NaturalReasoning, MMLU Pro, and SuperGPQA, we use the provided ground truth answers from each benchmark.
4.2 Profiling Harness
We develop an end-to-end, cross-platform profiling harness for inference workloads, ensuring reproducible results and support for new models, tasks, and hardware backends. It consists of three core components: distributed multi-GPU inference, response evaluation, and system-level telemetry collection. The harness supports three hardware backends: NVIDIA, macOS (Apple Silicon), and AMD. Given a dataset, model, and hardware backend, it orchestrates inference over all input queries, evaluates outputs (via exact match or LLM-as-a-judge), and records detailed telemetry—latency, throughput, time-to-first-token (TTFT), energy consumption, and more (see Table 8). Telemetry is collected via vendor APIs, synchronized at nanosecond resolution, and normalized (watts, joules, megabytes). For energy measurements, we follow standard practices (Samsi et al., 2023a; Fernandez et al., 2025; Wilkins et al., 2024): querying NVML’s energy counter for NVIDIA GPUs and using powermetrics with power integration over time for Apple Silicon. We differ in two key aspects: (1) we sample at 50ms intervals for higher temporal resolution (compared to 100ms (Samsi et al., 2023a) or 15 seconds (Fernandez et al., 2025)), and (2) for multi-GPU configurations, we aggregate energy counters from each GPU individually rather than extrapolating from a single device (Samsi et al., 2023a). On NVIDIA systems, we use NVML to retrieve per-device power, energy, memory usage, and temperature. On macOS systems, we extract GPU power data from powermetrics and compute energy via numerical integration. On AMD systems, we query ROCm SMI for power, temperature, and VRAM usage, again integrating power over time. Full details provided in App. B.1.
5 Intelligence Efficiency Study
Figure 2: Local Models Rival Cloud Models Across Diverse Benchmarks: Individual model performance scales with size, ranging from 28.8–72.5% for Qwen3-4B to 66.5–89.5% for Gemini 2.5 Pro. Local routing (best local LM per query) achieves 97.8%, 88.3%, 77.0%, and 92.4% on Wildchat, NaturalReasoning, SuperGPQA, and MMLU Pro respectively, surpassing cloud routing (100%, 82.9%, 66.5%, 87.4%) on three of four benchmarks.
We investigate whether recent advances in local LMs and local accelerators enable local inference to viably complement centralized cloud infrastructure by handling a substantial fraction of inference queries. Using our curated dataset, we examine three interconnected questions: 1) the extent to which current workloads can be handled locally (Section 5.1), 2) how intelligence efficiency has evolved from 2023-2025 (Section 5.2), 3) what gains query routing across local and cloud models can deliver in practice (Section 5.3). We study single-query inference (batch size = 1), to (1) isolate intrinsic model-accelerator efficiency from system-level serving optimizations and (2) follow standard local inference benchmarking practices (Hao et al., 2023).
5.1 Can Local Models and Accelerators Handle Current Inference Workloads?
Figure 2 shows query coverage (the percentage of dataset queries correctly answered) across individual local LMs, the best-of-local ensemble (routing to the best local LM for each query), and the best-of-cloud baseline (routing to the best frontier model).
Local LM coverage increases with scale and time. Across Wildchat, NaturalReasoning, SuperGPQA, and MMLU Pro, individual model coverage ranges from $49.6\%$ for Qwen3-4B, on average, to $71.4\%$ for gpt-oss-120b, with consistent improvements at each scale point: Qwen3-8B achieves $57.5\%$ and Qwen3-14B reaches $60.0\%$. Coverage has improved substantially from 2023 to 2025 (Figure 3): the best local LMs achieved a $32.2\%$ relative improvement on chat queries and a $50.1\%$ relative improvement on reasoning queries over this period. While improvements from 2023-2025 are relatively uniform across difficulty levels for chat tasks, reasoning tasks show markedly slower progress on the hardest problems (see App C). These results demonstrate that larger local LMs can handle progressively more queries without requiring cloud infrastructure, with the best individual local LM (gpt-oss-120b) successfully answering almost three-fourths of the single-turn chat and reasoning queries studied.
Model diversity substantially improves coverage. Routing queries to the most appropriate local LM rather than using a single model achieves $88.7\%$ overall coverage—a $28.8$ percentage point improvement over Qwen3-14B and $16.3$ percentage points over individual gpt-oss-120b performance, on average. This gap between individual models and best-of-local demonstrates that architectural, pretraining, and post-training diversity captures complementary capabilities: different models excel on different query types, and intelligent routing can exploit these complementary strengths.
Chat queries are more amenable to local processing than reasoning queries. Performance varies significantly by task type: Wildchat queries show substantially higher coverage across all models, while NaturalReasoning queries prove more challenging. The best local LM achieves $88.9\%$ coverage on Wildchat versus $64.9\%$ on NaturalReasoning —a $24.0$ percentage point gap. This aligns with findings that $77\%$ of real-world ChatGPT queries involve practical guidance, information seeking, or writing (Chatterji et al., 2025)—tasks well-suited to local models—while reasoning-intensive queries more often require frontier capabilities for tasks such as Architecture, Engineering, Life & Physical Science, and Mathematics (see Figure 7 in App. B). Nevertheless, based on our local routing setting for NaturalReasoning, SuperGPQA, and MMLU Pro (Figure 2), local LMs successfully handle over four-fifths of reasoning queries studied, suggesting significant opportunities for local inference even in technically demanding domains.
[…]
Evaluation on standardized benchmarks confirms local LM viability across task distributions. To validate that our findings generalize beyond naturalistic queries, we evaluate local model coverage on MMLU Pro (multi-domain knowledge) and SuperGPQA (graduate-level reasoning across 285 disciplines). The best individual local model (gpt-oss-120b) achieves $80.4\%$ on MMLU Pro and $51.5\%$ on SuperGPQA, while routing to the best local LM for each query achieves $93.4\%$ coverage on MMLU Pro and $83.6\%$ coverage on SuperGPQA, which is a $13.0$ and $32.1$ percentage point improvement, respectively. Notably, SuperGPQA’s comprehensive disciplinary coverage reveals that technical domains remain the primary challenge for local deployment: coverage exceeds $93.1\%$ for creative and humanities fields but drops to $60.1\%$ for specialized technical disciplines like Architecture & Engineering (Figure 7). This domain-specific analysis confirms that while local LMs can handle the majority of conversational and knowledge-recall tasks, complex reasoning in specialized fields still benefits from frontier model capabilities.
Figure 4: Increasing GPU Memory of Consumer Accelerators: Memory capacity (GB) for local accelerators. Over the past decade, local hardware has significantly closed the memory gap with cloud-grade accelerators, particularly since 2020, driven by advances in high bandwidth memory (HBM) components and unified memory architectures.
Local accelerator memory capacity is expanding rapidly. From 2012 to 2025, local accelerator memory has increased dramatically, with particularly rapid growth since 2020 (Figure 4). Local accelerators that offered 10-20 GB in 2020 now provide 128-512 GB through unified memory architectures like Apple Silicon, enabling models that previously required datacenter infrastructure to run efficiently on local hardware. This memory expansion has been the primary driver enabling local deployment of increasingly capable models: the jump from sub-20 GB to 200+ GB memory removes the key constraint that forced workloads to cloud infrastructure, allowing local accelerators to host the 8-20B active parameter models that now handle the majority of inference queries.
5.2 How Intelligence Efficient is Local Inference?
Intelligence efficiency is improving over time. Table 2 tracks the evolution of local LM capabilities from 2023 to 2025, measuring the best available local LM ($\leq 20B$ active parameters) paired with state-of-the-art accelerators each year. On our curated dataset of chat and reasoning queries, accuracy per watt has improved $5.3\times$ over this two-year period: in 2023, Mixtral-8x7B-v0.1 on NVIDIA Quadro RTX 6000 achieved $7.92\times 10^{-4}$ accuracy per watt; by 2024, Llama-3.1-8B-Instruct on NVIDIA RTX 6000 Ada reached $1.80\times 10^{-3}$ (a $2.27\times$ year-over-year gain); and in 2025, gpt-oss-120b on Apple M4 Max achieved $4.18\times 10^{-3}$ (a $2.32\times$ gain). Notably, local LM coverage on single-turn chat and reasoning queries has increased in lockstep with efficiency gains: from $23.2\%$ in 2023 to $48.7\%$ in 2024 to $71.3\%$ in 2025. This progression reflects compounding improvements in both model architectures, which achieve higher accuracy through advances in pretraining (Chowdhery et al., 2023; Hoffmann et al., 2022; OpenAI, 2023; DeepSeek-AI, 2024), post-training (Bai et al., 2022; Shao et al., 2024b; DeepSeek-AI, 2025), and parameter utilization via mixture-of-experts (MoE) architectures (Shazeer et al., 2017; DeepSeek-AI, 2024), and hardware accelerators, which deliver more compute (FLOPs) and memory per watt (NVIDIA, 2021; 2024).
Figure 5 provides a complementary view by examining accuracy per watt trends across model releases between 2023 and 2025, showing consistent efficiency improvements across multiple model families: (Llama, Phi, Gemma, Mistral, Falcon, DeepSeek, Qwen, and gpt-oss). The consistent progression across multiple efficiency metrics—including accuracy per joule and perplexity-based measurements presented in Figure 10 (App. C.2)—demonstrates that intelligence efficiency is improving rapidly and predictably. As both model architectures and hardware accelerators continue to advance, the viability of local inference will expand further, enabling an increasingly large fraction of queries to be served efficiently outside centralized cloud infrastructure.
2023 2024 2025
SOTA
Local Model
Mixtral-8x7B-v0.1 Llama-3.1-8B-Instruct GPT-OSS-120B
SOTA
Accelerator
NVIDIA Quadro
RTX 6000
NVIDIA RTX
6000 Ada
Apple
M4 Max
Success Rate
23.2% 48.7% 71.3%
Intelligence
per Watt
$7.92\times 10^{-4}$ $1.80\times 10^{-3}$ $4.18\times 10^{-3}$
YoY Efficiency
Gain
2.27× 2.32×
Table 2: Increase in Intelligence per Watt for Local LMs: Accuracy per watt has improved over $5\times$ in two years, driven by advances in both model architectures (from Mixtral-8x7B to GPT-OSS-120B) and accelerator hardware (from NVIDIA Quadro RTX 6000 to Apple M4 Max).
Figure 5: Increase in Intelligence per Joule for Local LMs and Accelerators: Efficiency improved $18.0\times$ over 16 months, decomposed into $3.1\times$ from better local LMs and $5.9\times$ from better local accelerators.
Local accelerator efficiency has room for improvement: While local accelerators enable deployment of capable models outside data centers, cloud-grade hardware maintains a substantial efficiency advantage for the same workloads. We evaluate this gap using two complementary metrics: intelligence per watt (instantaneous power efficiency) and intelligence per joule (end-to-end energy efficiency per query).
Table 3 compares intelligence per watt across Qwen3 models (4B to 32B active parameters) on Apple M4 Max (local) versus NVIDIA B200 and SambaNova SN40L (cloud). The B200 achieves $1.40\times$ higher intelligence per watt than the M4 Max across all model sizes, while the SN40L achieves $1.78\times$ higher efficiency on Qwen3-32B ($3.51\times 10^{-3}$ versus $1.97\times 10^{-3}$ intelligence per watt).
Table 4 extends this analysis to intelligence per joule, which accounts for both power consumption and generation latency across Qwen3 and GPT-OSS model variants. Here the efficiency gaps widen substantially: the B200 achieves $1.6\times$ to $2.3\times$ higher intelligence per joule than the M4 Max, while the SN40L achieves $6.5\times$ to $7.4\times$ higher efficiency. The larger gap in per-joule metrics reflects that cloud accelerators not only consume less power per unit of accuracy, but also complete queries faster, compounding their energy advantage.
These efficiency gaps stem from specialized hardware optimizations in enterprise-grade accelerators: cloud accelerators like the B200 and SN40L employ purpose-built components—high-bandwidth memory (HBM3e), dedicated tensor processing units, and optimized memory hierarchies—that maximize throughput per watt, whereas local accelerators use unified memory architectures that balance diverse workloads under thermal and power constraints. Notably, all measurements use batch size 1, indicating these advantages persist even in single-query settings and revealing substantial headroom for future local accelerator designs to close this gap through specialized on-device AI components. However, this efficiency disadvantage is offset by complementary system-level benefits: local deployment avoids datacenter infrastructure costs and enables $88.7\%$ of queries that local models can handle to avoid cloud compute entirely, yielding $60{-}80\%$ resource reductions through intelligent routing (Section 5.3).
Qwen3-4B Qwen3-8B Qwen3-14B Qwen3-32B
Success Rate
49.3% 57.5% 59.5% 69.5%
Apple M4 Max
Intelligence
per Watt
$1.40\times 10^{-3}$ $1.63\times 10^{-3}$ $1.69\times 10^{-3}$ $1.97\times 10^{-3}$
NVIDIA B200
Intelligence
per Watt
$1.95\times 10^{-3}$ $2.27\times 10^{-3}$ $2.35\times 10^{-3}$ $2.75\times 10^{-3}$
SambaNova SN40L
Intelligence
per Watt
$3.51\times 10^{-3}$
Table 3: Local accelerators demonstrate lower power efficiency than cloud accelerators: When running the same Qwen3 models, the Apple M4 Max (local) attains $1.40\times$ lower intelligence per watt compared to the NVIDIA B200 (cloud) and SambaNova SN40L (cloud), highlighting the efficiency advantage of purpose-built cloud accelerators over local accelerators.
Qwen3-8B Qwen3-32B GPT-OSS-20B GPT-OSS-120B
Apple M4 Max
Intelligence
per Joule
$3.80\times 10^{-5}$ $3.51\times 10^{-5}$ $4.38\times 10^{-5}$ $4.23\times 10^{-5}$
NVIDIA B200
Intelligence
per Joule
$8.71\times 10^{-5}$ $5.91\times 10^{-5}$ $7.34\times 10^{-5}$ $6.78\times 10^{-5}$
SambaNova SN40L
Intelligence
per Joule
$2.27\times 10^{-4}$ $3.12\times 10^{-4}$
Table 4: Cloud accelerators demonstrate superior energy efficiency across all models: The NVIDIA B200 (cloud) achieves $1.6\times$ to $2.3\times$ higher intelligence per joule than the Apple M4 Max (local), while the SambaNova SN40L (cloud) achieves $6.5\times$ to $7.4\times$ higher efficiency. These results highlight the substantial energy efficiency advantage of purpose-built cloud accelerators over local hardware across Qwen3 and GPT-OSS model variants.
Figure 6: Energy, Compute, and Capital Gains from Model Routing. Cumulative resource consumption over 24 hours and 80.2M LLM queries (Wang et al., 2025). Using our local-cloud router between 44 small LMs on Apple M4 Max and Qwen3-235B on an H200 yields substantial savings at various routing accuracies. For the 80% accurate router, we observe: 64.3% in energy savings, 61.8% in compute, and 59.0% in cost compared to naively routing every query to Qwen3-235B, capturing a majority of the gains achievable by the theoretical best-case (Oracle Router).
5.3 What Efficiency Gains Can Effective Query Routing Deliver?
We simulate a hybrid local-cloud system serving 80.280.2M queries over a 24-hour period—representative of realistic daily inference workloads (Wang et al., 2025). Queries are routed between four small local LMs (Qwen3-4B, Qwen3-8B, Qwen3-14B, gpt-oss-20b) running on Apple M4 Max devices and a frontier model (Qwen3-235B) running on NVIDIA H200 infrastructure. Figure 6 shows cumulative energy consumption (left), compute usage (middle), and cost (right) over the 24-hour period for five routing strategies: routing all queries to the largest model (baseline), oracle routing (perfect assignment), and two realistic routers with $60\%$ and $80\%$ routing accuracy. For the realistic routers, when the system fails to correctly identify the smallest capable local model, queries default to the cloud model (Qwen3-235B) to ensure correctness.
Oracle routing establishes theoretical upper bounds. Assuming perfect query-to-model assignment, oracle routing reduces energy consumption by $80.4\%$, compute by $77.3\%$, and cost by $73.8\%$ versus cloud-only deployment to the largest model (Figure 6). These dramatic savings stem from assigning the $80.7\%$ of queries that local models can handle correctly to significantly more efficient hardware, while reserving expensive frontier model compute for the remaining $19.3\%$ of queries that truly require it.
Practical routers achieve substantial gains without perfect accuracy. Crucially, routing systems need not achieve perfect accuracy to realize significant efficiency improvements. A router with $80\%$ accuracy—a realistic target for modern routing systems—captures approximately $80\%$ of oracle gains, achieving $64.3\%$ energy reduction, $61.8\%$ compute reduction, and $59.0\%$ cost reduction compared to the cloud-only baseline (Figure 6). Even a more conservative $60\%$ accurate router delivers $48.4\%$ energy savings, $46.7\%$ compute savings, and $44.5\%$ cost savings. Importantly, these efficiency gains come with minimal accuracy degradation: routing at $80\%$ accuracy maintains task performance as misrouted queries fall back to the frontier model. These savings scale linearly with query volume: at current platform-scale inference demand (billions of queries daily), intelligent routing could yield annual energy savings measured in terawatt-hours.
6 Extended Experiments
Finally, we include a set of extended experiments in the Appendix, including:
  • How does model precision affect task performance and intelligence efficiency (Appendix C.4)?
  • How do SOTA open-source LMs compare to the SOTA closed-source LMs on Chat and Reasoning Queries (Appendix C.5)?
  • How does performance on chat and reasoning queries connect to U.S. GDP (Appendix C.6)?
7 Conclusion
Centralized cloud infrastructure faces mounting pressure from exponential growth in LLM inference demand. Advances in local models and accelerators create an opportunity to redistribute workloads, but evaluating this transition requires quantifying both model capability and hardware efficiency. We introduce intelligence per watt as a unified metric for measuring local inference viability and conduct a large-scale empirical study across 20+ models, 8 hardware accelerators, and 1M real-world queries spanning 2023-2025. Our work makes three contributions. First, we demonstrate that local LMs can successfully handle $88.7\%$ of single-turn chat and reasoning queries, with coverage varying by domain—exceeding $90\%$ for creative tasks but falling to $68\%$ for technical fields. Second, we show that intelligence per watt has improved $5.3\times$ from 2023-2025 through compounding advances in both model architectures ($3.1\times$) and hardware accelerators ($1.7\times$), with locally-serviceable query coverage increasing from $23.2\%$ to $71.3\%$. Third, we demonstrate that treating local and cloud infrastructure as complementary resources and optimally routing queries across them (based on the smallest model capable of answering the queries), achieves $60{-}80\%$ reductions in energy, compute, and cost through intelligent query routing, even with realistic ($80\%$) routing accuracy, while maintaining answer quality. These findings establish that local inference can meaningfully redistribute demand from centralized infrastructure for a substantial and growing subset of queries. As model architectures and hardware accelerators continue to advance, intelligence per watt will remain the critical metric for tracking this transition. We release our profiling harness to enable systematic efficiency benchmarking as the local inference ecosystem evolves.
8 Acknowledgements
We thank Simran Arora, Bradley Brown, Mayee Chen, Owen Dugan, Neel Guha, Simon Guo, Jordan Juravsky, Jerry Liu, Benjamin Spector, and Shayan Talaei for their constructive feedback during the composition of the paper. We would also like to thank our collaborators at the Stanford Artificial Intelligence Laboratory (SAIL) and TogetherAI.
We gratefully acknowledge the support of NIH under No. U54EB020405 (Mobilize); NSF under Nos. CCF2247015 (Hardware-Aware), CCF1763315 (Beyond Sparsity), CCF1563078 (Volume to Velocity), and 1937301 (RTML); US DEVCOM ARL under Nos. W911NF-23-2-0184 (Long-context) and W911NF-21-2-0251 (Interactive Human-AI Teaming); ONR under No. N000142312633 (Deep Signal Processing); Stanford HAI under No. 247183; Google DeepMind; Google Research; Google Cloud; NXP; Xilinx; LETI-CEA; Intel; IBM; Microsoft; NEC; Toshiba; TSMC; ARM; Hitachi; BASF; Accenture; Ericsson; Qualcomm; Analog Devices; Salesforce; Total; the Laude Institute; Prime Intellect; Together; Anthropic; the HAI-GCP Cloud Credits for Research program; the Stanford Data Science Initiative (SDSI); members of the Stanford DAWN project: Meta, Google, and VMWare; and members of the Stanford SEAMS project: IBM and Felicis. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views, policies, or endorsements, either expressed or implied, of NIH, ONR, or the U.S. Government.
Appendix A Extended Related Works
Below we provide an extended treatment of related works.
LLM Routing
A central challenge in local-cloud routing systems is determining which model should handle a given query so as to maximize efficiency. Prior work spans a broad design space, but much of it can be organized around two families of approaches: embedding-based routers (Zhang, 2025a; Somerstep et al., 2025; Chen et al., 2024) and generative/decoder-based routers (Ong et al., 2025). Embedding-based methods rely on encoding queries (and sometimes models) into a vector space and then applying similarity search or lightweight classification. Early work largely adopted binary routing, where queries are directed between just two models. For example, RouteLLM (Ong et al., 2024) demonstrated that simple supervised classification can yield up to 85% cost reduction while maintaining GPT-4–level performance, but this setting was restricted to two-model scenarios. More recent systems generalize routing to multi-model settings: ensemble-style methods such as FrugalGPT (Chen et al., 2023), RouterDC (Chen et al., 2024), and Avengers Pro (Zhang, 2025a; b) show that intelligently combining smaller models can approximate or even surpass larger frontier LMs. Decoder-based methods leverage a small language model to directly generate the routing decision. Causal LLM Routing, suggests that incorporating richer query–model interaction signals via generative modeling or cross-attention can yield more robust routing than static embeddings (Chen et al., 2024). In this work, we are inspired by these novel approaches to routing, and evaluate their performance in the local-cloud routing setup.
LLM Routing Benchmarks
Recent work has explored benchmarks for LLM query routing, primarily targeting cost–quality tradeoffs across multiple models. RouterBench (Hu et al., 2024) provides a comprehensive suite of curated academic tasks ( 405K samples) to evaluate routing policies along cost–quality Pareto frontiers. RouteLLM (Ong et al., 2024) introduces a preference-trained routing framework evaluated on academic benchmarks like MMLU and MT-Bench, with a focus on achieving quality under token cost constraints, though it remains limited to token-level metrics. RouterEval (Huang et al., 2025) emphasizes model selection accuracy at scale, compiling over 200M performance records across 8.5K models and 12 benchmarks to study generalization, yet lacks coverage of real-world queries. In contrast, our curated dataset targets routing under naturalistic conditions, leveraging 1M real user queries from Wildchat and NaturalReasoning. It uniquely supports the exploration of local-cloud routing tradeoffs beyond just cost and quality, to metrics such as latency, energy, memory, throughput, and more — generated on local accelerators and enterprise-grade accelerators. Moreover, in contrast to existing benchmarks—which provide stale performance records limited to models released prior to July 2024—our curated dataset evaluates several state-of-the-art models, including Qwen3 (Team, 2025) and GPT-OSS (Agarwal et al., 2025), all released after May 2025. To support ongoing benchmarking, we release our efficiency profiling harness, a hardware-agnostic toolkit for generating fresh telemetry and evaluation records as new models become available.
Local–Cloud Inference Systems
Beyond model selection, recent work explores collaborative inference protocols that split generation between local and cloud models. Minions (Narayan et al., 2025) proposes a two-stage protocol where a small on-device LM handles lightweight processing and a frontier LM performs high-level reasoning, with an extended version introducing task decomposition and aggregation for improved quality. Such collaborative schemes offer large energy and cost savings but require careful protocol design to avoid performance loss. A parallel line of work centers on speculative decoding, where a small draft model generates candidate continuations that are verified or refined by a larger target LM (Miao et al., 2023; Xu et al., 2025). These approaches primarily target latency and throughput, particularly in constrained hardware settings, and typically assume that generation will ultimately invoke a large LM. Other hybrid protocols like SLED (Li et al., 2025) and HAT (Xie et al., 2025) introduce edge–cloud model partitioning with intermediate state exchange to balance device limitations with quality needs. While these systems explore fine-grained collaboration at the token or layer level, our work investigates the limitations of a coarser-grained alternative: query-level routing across multiple small and large LMs, where we measure not only accuracy and cost, but also latency, memory, and energy across diverse hardware accelerators.
Efficient AI
We are inspired by work on ”Green AI” which proposes treating energy as a first-class metric alongside accuracy and cost, with calls for standardized reporting and tooling for reproducible accounting of power use and emissions during training and inference (Schwartz et al., 2020; Strubell et al., 2019; Patterson et al., 2021; Henderson et al., 2020; Anthony et al., 2020; CodeCarbon Contributors, 2024; Oviedo et al., 2025). Complementary to our focus on local–cloud routing, cost and efficiency-driven model selection strategies such as FrugalML and FrugalGPT for API and model cascades, and CALM for token-wise early exit, dynamically allocate workloads to cheaper or smaller models while preserving quality (Chen et al., 2020; 2023; Schuster et al., 2022). On-device and edge studies demonstrate algorithm–hardware co-design for lower latency and energy consumption, exemplified by EdgeBERT’s optimizations and PowerInfer’s efficient LLM serving on commodity GPUs (Tambe et al., 2021; Song et al., 2024). Finally, hardware-aware benchmarking efforts such as ”From Words to Watts” and MLPerf Power quantify inference energy across accelerators and standardize power measurement protocols (Samsi et al., 2023b; Tschand et al., 2025).
Appendix B Dataset and Profiling Harness
In this section, we provide additional details on our dataset curation for our study of hybrid local-cloud LM systems.
B.1 Dataset Curation
Here, we provide additional details on the Anthropic Economic Index (Handa et al., 2025) categories used as labels (see Table 5), the hardware platforms profiled, and the metrics recorded in our curated dataset.
Life, physical, and social science Computer and mathematical
Architecture and engineering Education instruction and library
Installation, maintenance, and repair Business and financial operations
Legal services Transportation and material moving
Arts, design, sports, entertainment, and media Production services
Farming, fishing, and forestry Healthcare support
Food preparation and serving related Healthcare practitioners and technical
Community and social service Sales and related
Office and administrative support General management
Protective service Building grounds cleaning and maintenance
Construction and extraction Personal care and service
Table 5: Anthropic Economic Index Categories (Handa et al., 2025). This taxonomy categorizes occupations into 22 standardized economic domains, adapted from U.S. Bureau of Labor Statistics frameworks. It is designed to support AI impact analysis by aligning labor categories with distinct task structures.
Query Curation
When sourcing queries from the Wildchat and NaturalReasoning datasets, we apply robust data cleaning and filtering to ensure the quality and consistency of the sampled queries. For NaturalReasoning, we filter out all queries that don’t contain ground truth answers. For Wildchat, we eliminate non-English entries to maintain linguistic uniformity across the dataset. Queries that are malformed, nonsensical, or otherwise unintelligible (as determined by an LLM judge—i.e., GPT-4o-mini) are discarded to prevent noise. Additionally, duplicate queries are removed to reduce redundancy and avoid overrepresentation of specific prompts. Finally, we filter out excessively long queries that exceed a 32,000-character limit.
Dataset Statistics
Table 6 reveals significant differences in how the two datasets are distributed across domains. Wildchat is dominated by “Arts, design, sports, entertainment, and media” queries (47.1%), followed by “Computer and mathematical” (18.1%), while NaturalReasoning is primarily composed of “Life, physical, and social science” (36.0%) and “Computer and mathematical” (34.8%) queries. We use gpt-4o-mini to bucket each query into its economic categorization using the prompt below.
1You are a query categorizer. Your task is to categorize the following user query into one of the predefined categories based on the job/occupation domain it relates to most closely.
2
3 Query: "{query}"
4
5 Available Categories:
6 - Office and administrative support
7 - Transportation and material moving
8 - Sales and related
9 - Food preparation and serving related
10 - General management
11 - Business and financial operations
12 - Healthcare practitioners and technical
13 - Production services
14 - Education instruction and library
15 - Healthcare support
16 - Construction and extraction
17 - Installation, maintenance, and repair
18 - Computer and mathematical
19 - Building grounds cleaning and maintenance
20 - Protective service
21 - Personal care and service
22 - Architecture and engineering
23 - Community and social service
24 - Arts, design, sports, entertainment, and media
25 - Life, physical, and social science
26 - Legal services
27 - Farming, fishing, and forestry
28 - None
29
30 Instructions:
31 1. Read the query carefully
32 2. Determine which job/occupation category the query relates to most closely
33 3. If the query doesnt clearly relate to any specific occupation category, use "None"
34 4. Respond with ONLY the category name, exactly as listed above
35
36 Category:
Solvability rates vary dramatically by domain and dataset type, where a query’s solvability is defined as its ability to be answered correctly by any of the available local LMs (e.g. Qwen models or GPT OSS). Wildchat queries show consistently high solvability across most domains (generally $94\%$), with particularly strong performance in creative and social domains. In contrast, NaturalReasoning exhibits more variable solvability, with technical domains like “Architecture and engineering” showing only 41.5% solvability compared to 99.4% for the same domain in Wildchat. This disparity reflects the complexity difference between open-ended chat queries and analytical reasoning tasks, supporting our findings that chat queries are more amenable to local model routing than reasoning-intensive queries.
Figure 7: Local Win/Tie-Rate vs. Cloud LMs by Domain. Stacked bars show the fraction of single-turn chat and reasoning queries handled by local LMs (<20B<20B active parameters; blue) versus those routed to frontier models in the cloud (red), computed per economic index domain (Appel et al., 2025)
Domain
WC
Count
WC
WC
Solv %
NR
Count
NR
NR
Solv %
Computer and mathematical 90,662 18.1 99.5 174,242 34.8 67.3
Arts, design, sports, entertainment,
and media
235,658 47.1 98.7 2,648 0.5 52.9
Life, physical, and
social science
28,079 5.6 98.8 180,065 36.0 60.5
None 49,014 9.8 97.3 79,752 16.0 65.6
Education instruction and library 23,196 4.6 97.2 13,864 2.8 80.4
Architecture and engineering 5,782 1.2 98.9 28,762 5.8 40.8
Business and financial operations 19,628 3.9 97.8 8,779 1.8 55.3
Healthcare practitioners
and technical
8,905 1.8 98.1 1,851 0.4 66.3
Office and administrative
support
6,959 1.4 91.6 25 0.0 48.0
Legal services 5,208 1.0 98.6 1,349 0.3 69.0
Community and social service 6,125 1.2 97.0 404 0.1 76.2
Transportation and
material moving
1,890 0.4 95.1 3,914 0.8 53.0
Sales and related 4,689 0.9 97.8 218 0.0 67.4
Food preparation and
serving related
3,308 0.7 98.3 507 0.1 61.9
General management 3,364 0.7 96.7 340 0.1 67.9
Installation, maintenance,
and repair
954 0.2 97.1 2,037 0.4 56.2
Farming, fishing, and forestry 1,677 0.3 99.5 411 0.1 65.5
Protective service 1,315 0.3 97.9 237 0.0 56.5
Construction and extraction 991 0.2 97.2 147 0.0 60.5
Healthcare support 1,112 0.2 97.5 12 0.0 100.0
Production services 546 0.1 100.0 334 0.1 65.3
Personal care and service 648 0.1 92.9 19 0.0 0.0
Building grounds cleaning
and maintenance
278 0.1 100.0 70 0.0 72.9
TOTAL 500K 100.0 98.4 500K 100.0 63.0
Table 6: Dataset Domain Composition and LM Coverage ($\leq 20B$ Active Parameter Models). Comparison of domain distribution and model solvability rates across Wildchat (WC) and NaturalReasoning (NR) datasets. Solvability indicates the percentage of problems that can be solved correctly by at least one model with $\leq 20B$ active parameters.
Category Wildchat MMLU Pro SuperGPQA Average
Computer and mathematical 93.4% 90.6% 72.8% 85.6%
Life, physical, and social science 91.1% 84.7% 50.4% 75.4%
Sales and related 86.8% 74.2% 64.3% 75.1%
Business and financial operations 89.3% 82.9% 52.5% 74.9%
Production services 89.7% 85.7% 48.8% 74.7%
Office and administrative support 88.5% 83.3% 44.8% 72.2%
Healthcare practitioners and technical 88.9% 78.8% 48.0% 71.9%
Installation, maintenance, and repair 90.4% 80.0% 44.4% 71.6%
Architecture and engineering 90.0% 73.2% 51.6% 71.6%
Protective service 88.6% 75.0% 47.6% 70.4%
Education instruction and library 90.6% 77.4% 43.2% 70.4%
Farming, fishing, and forestry 87.2% 77.8% 45.9% 70.3%
None 86.5% 77.4% 42.4% 68.8%
General management 88.2% 76.3% 41.6% 68.7%
Transportation and material moving 91.1% 66.7% 44.9% 67.5%
Construction and extraction 91.0% 66.7% 41.4% 66.4%
Food preparation and serving related 86.1% 82.6% 30.3% 66.3%
Community and social service 87.4% 64.1% 45.5% 65.7%
Arts, design, sports, entertainment, and media 84.8% 75.5% 36.2% 65.5%
Building grounds cleaning and maintenance 90.1% 71.4% 30.8% 64.1%
Legal services 87.1% 61.5% 43.6% 64.1%
Healthcare support 86.5% 60.0% 38.9% 61.8%
Personal care and service 89.4% 50.0% 25.0% 54.8%
Table 7: GPT-OSS-120B Performance across Datasets. Performance metrics across Wildchat, NaturalReasoning, MMLU Pro, and SuperGPQA benchmarks, organized by Anthropic Economic Index categories (Handa et al., 2025).
Metrics
We detail all the metrics collected via our profiling harness in Table 8. For our correctness evaluations on Wildchat, we use an LLM-as-a-judge approach to evaluate model generated answers against a ground truth answer from Qwen3-235B. For our correctness evaluations of NaturalReasoning, we use another LLM-judge prompt, but compare against ground truth answers provided in the original dataset. We provide both LLM-judge prompts below. The LLM used for each respective evaluation is GPT-4o. For SuperGPQA and MMLU Pro, we simply compare the multiple choice answer selected in the response to the multiple choice answer of the reference response.
Wildchat LLM-judge Prompt
1
2You are an impartial judge evaluating the quality of two AI-assistant replies to the same user prompt.
3
4Step 1: Generate your own answer
5Write the response *you* would give to the user. Keep it separate from later analysis.
6
7Step 2: Decide the query type
8Classify the user prompt as either
9- **Subjective / open-ended** (creative writing, opinion, advice, brainstorming)
10- **Objective / technical** (code, math, logical derivations with a single correct outcome)
11If uncertain, default to "Subjective".
12
13Step 3 - Score each assistant with the correct rubric
14
15| Query type | Criteria |
16|------------|----------|
17| Subjective / open-ended | 1. Correctness / factual soundness 2. Helpfulness 3. Relevance 4. Conciseness 5. Creativity & novelty |
18| Objective / technical | 1. Correctness only |
19
20When using the multi-criteria rubric, note strengths and weaknesses for **each** dimension.
21When using the single-criterion rubric, focus exclusively on factual / functional accuracy and ignore style or flair.
22
23Step 4: Compare & justify
24Explain which assistant is better and why, correcting any mistakes you find. Highlight missing but important details. **Be concise.**
25
26Step 5: Verdict
271. Assistant A is significantly better: [[A>>B]]
282. Assistant A is slightly better: [[A>B]]
293. Tie, Assistant A is equal: [[A=B]]
304. Assistant B is slightly better: [[B>A]]
315. Assistant B is significantly better: [[B>>A]]
32
33Choose exactly one token from: ‘[[A>>B]]‘, ‘[[A>B]]‘, ‘[[A=B]]‘, ‘[[B>A]]‘, ‘[[B>>A]]‘.
34
35---
36
37### Output format (strict)
38Return **only** a JSON object that matches the provided schema:
NaturalReasoning LLM-judge Prompt
1You are evaluating a response to a scientific/technical question against a reference answer.
2
3Your task is to determine if the response is factually correct and complete compared to the reference.
4
5Consider:
61. Scientific accuracy of facts and concepts
72. Mathematical correctness (if applicable)
83. Completeness of the answer
94. Technical precision
10
11Question: {question}
12
13Response: {response}
14
15Reference Answer: {reference}
16
17Return ONLY True if the response is correct and complete, False otherwise.
Metric Description
flops_per_request FLOPs per query.
macs_per_request MACs per query; proxy for compute.
per_query_joules Energy per query (J).
total_joules Total energy across queries.
per_token_ms Latency per token (ms).
throughput_tokens_per_sec Token output rate (toks/s).
time_to_first_token_seconds Time to first token (s).
total_query_seconds Total time per query (s).
cpu_mb.avg / max / median / min CPU memory usage (MB).
gpu_mb.avg / max / median / min GPU memory usage (MB).
initialization_duration_seconds Model load time (s).
batch_size Query batch size.
gpu_memory_utilization GPU memory use (0–1).
max_model_len Max token length allowed.
max_num_batched_tokens Max batch token count.
max_output_tokens Max output tokens.
num_workers Number of threads.
temperature Sampling temperature.
top_k Top-k cutoff.
top_p Top-p (nucleus) threshold.
warmup_steps Warm-up steps.
per_query_watts.avg / max / median / min GPU power draw per query (W).
total_watts.avg / max / median / min Session GPU power draw (W).
cpu_count CPU core count.
cpu_brand CPU model.
host_name Machine hostname.
os_name / os_version / kernel_version OS and kernel info.
temperature.avg / max / median / min Device temperature (°C).
input Input tokens per query.
output Output tokens per query.
Table 8: Dataset Metrics. Summary of compute, latency, memory, and energy profiling metrics.
Hardware Backends
Details regarding profiled hardware can be found in Table 9.
Data Generation Procedure
We generate model outputs using consistent decoding settings across all tasks: temperature == 0.6, top-p == 0.95, top-k == 20, min-p == 0.0, and a 32768-token output limit. For NaturalReasoning, SuperGPQA and gpqa queries, we enable deliberative prompting (use thinking == True); for Wildchat, we disable it. For Qwen models, we apply a repetition penalty of 1.1 and length penalty of 1.0.
Hardware Memory Bandwidth Power
NVIDIA A100 40 GB SXM4 (Ampere) 40 GB HBM2 1,555 GB/s 400 W TDP
NVIDIA H200 SXM (Hopper) 141 GB HBM3e 4.8 TB/s Up to 700 W TDP
NVIDIA B200 (Blackwell) 192 GB HBM3e 8 TB/s 1000 W TDP
NVIDIA GH200 (Grace Hopper) 144 GB HBM3e + 624 GB LPDDR5X 4.8 TB/s (GPU) 1000 W TDP
NVIDIA Quadro RTX 6000 (Turing) 24 GB GDDR6 672 GB/s 295 W TDP
NVIDIA RTX 6000 Ada Generation 48 GB GDDR6 960 GB/s 300 W TDP
AMD Instinct MI300X (CDNA 3, OAM) 192 GB HBM3 5.3 TB/s (peak) 750 W TBP
Apple Mac Studio (M4 Max) 128 GB unified 546 GB/s 480 W (continuous)
SambaNova SN40L RDU 64 GB HBM3 + 1.5 TB DDR 2 TB/s (HBM) 600 W TDP
Table 9: Accelerator Details. Memory, bandwidth, and power specifications of evaluated accelerators and systems.
Telemetry Collection
We collected telemetry by instrumenting host-level samplers that interface directly with vendor-supported system APIs on each platform. Data were obtained from NVML on NVIDIA-equipped hosts, from the powermetrics facility on macOS, and from ROCm SMI on AMD-equipped hosts. Each sampler queried the respective system interface to obtain GPU- and system-level measurements and produced synchronized records suitable for downstream quantitative analysis.
On NVIDIA systems, we interface directly with NVML and enumerate all visible GPUs. For each device, we query instantaneous power as reported by the driver, read cumulative energy from the on-device counter, obtain GPU temperature from the hardware sensor, and retrieve memory usage from the device’s memory interface. Units are normalized (e.g., milliwatts and millijoules mapped to watts and joules; bytes to megabytes). In multi-GPU hosts, power and memory are summed across devices and temperature is averaged to yield a single aggregate view. Each record also includes host memory usage from OS counters, a nanosecond timestamp, and device identity and backend provenance.
On macOS, we execute powermetrics with elevated privileges and ingest its continuous plist stream. Each plist frame is parsed to extract the GPU power value exposed by the system (Apple Silicon: processor_power.actual; Intel: processor.combined_power), which is normalized to watts. Energy (joules) is obtained by numerically integrating the power signal over successive frames using the measured inter-frame wall-clock interval. In parallel, system memory usage is sampled from OS counters. Every observation is timestamped and annotated with Apple device identity and an explicit powermetrics backend tag.
On AMD systems, we use ROCm SMI to query current GPU power (watts), read temperature from junction or edge sensors (°C), and obtain VRAM usage from the device memory interface (bytes to megabytes). Energy (joules) is computed by integrating the power signal over time using consecutive sampling intervals. System memory usage is read from OS counters. In multi-GPU machines, the primary device under observation is explicitly selected (GPU index 0 in our setup), and all records carry precise timestamps together with device identity and backend metadata.
Cost Savings Compute Savings Energy Savings
Size Threshold ($\leq$)
Qwen
+ GPT-OSS
Qwen
Qwen
+ GPT-OSS
Qwen
Qwen
+ GPT-OSS
Qwen
4B 65.2% 65.2% 65.1% 65.1% 63.5% 63.5%
8B 80.8% 80.8% 83.1% 83.1% 79.6% 79.6%
14B 89.0% 89.0% 93.0% 93.0% 87.0% 87.0%
20B 90.5% 97.4% 89.4%
32B 91.3% 91.9% 97.4% 92.8% 90.4% 90.5%
120B 91.1% 97.7% 90.7%
Table 10: Cost, Compute, and Energy Savings from Local-Cloud Routing on WildChat: Savings across different resources while maintaining task accuracy of SOTA open-source cloud model (i.e. Qwen3 235B-A22B).
Cost Savings Compute Savings Energy Savings
Size Threshold ($\leq$)
Qwen
+ GPT-OSS
Qwen
Qwen
+ GPT-OSS
Qwen
Qwen
+ GPT-OSS
Qwen
4B 52.9% 52.9% 54.5% 54.5% 46.3% 46.3%
8B 60.5% 60.5% 62.5% 62.5% 54.0% 54.0%
14B 68.7% 68.7% 70.1% 70.1% 62.5% 62.5%
20B 73.3% 72.2% 67.8%
32B 76.9% 75.9% 75.1% 75.1% 72.4% 71.6%
120B 86.7% 86.0% 86.0% 85.9%
Table 11: Cost, Compute, and Energy Savings from Local-Cloud Routing on NaturalReasoning: Savings across different resources while maintaining task accuracy of SOTA open-source cloud model (i.e. Qwen3 235B-A22B).
Appendix C Local-Cloud Experiments
Figure 8: Chat Task Performance by Difficulty Level and Year. Model success rates across four difficulty levels and three model generations (2023, 2024, 2025). The data reveals dramatic progress across all difficulty levels, with 2023 models achieving 28.79% overall success rising to 98.12% by 2025. Notably, Levels 1-3 approach near-perfect performance (98-99%), while Level 4 shows the largest relative improvement (+210.4% per year) despite starting from the lowest baseline (17.77%).
Figure 9: Reasoning Task Performance by Difficulty Level and Year. Model success rates on across five difficulty levels and three model generations. The benchmark shows a three-tier saturation pattern: near-complete (98-99% on Levels 1-2), approaching saturation (85-92% on Levels 3-4), and wide-open frontier (51% on Level 5).
C.1 How has local LM task coverage changed over different “difficulty” slices of the data
Using labels for query difficulty we quantify the rate of improvement of local LMs across task difficulty slices. We label each query by the minimum model size (in parameters) required to solve it when considering the SOTA LMs as of August 2025, categorizing queries into five difficulty levels: level 1 ($\leq$ 4B params), level 2 ($\leq$ 8B params), level 3 ($\leq$ 20B params), level 4 ($\leq$ 235B params), and level 5 (unsolvable).
For chat tasks (see Figure 8), we observe near-universal performance gains across all difficulty levels, with 2025 models achieving 98-99% success on levels 1-3 and 92.6% on level 4. Absolute improvements range from +55.4 percentage points (pp) for level 1 to +76.4 pp for level 3, indicating relatively uniform capability gains.
For reasoning tasks (see Figure 9, the pattern differs substantially. While levels 1-3 show strong improvements (+24.0, +37.8, and +53.9 pp respectively), levels 4 and 5 exhibit markedly slower progress. Level 4 improves by only +23.8 pp (7.93% to 31.72%), and level 5 remains largely unsolved with just +1.5 pp improvement (3.27% to 4.72%). This suggests that while local models have rapidly closed the gap on moderately difficult reasoning tasks, the hardest reasoning problems—those requiring either massive scale or capabilities beyond current architectures—remain a significant frontier. The presence of 134 level 5 problems (16.5% of the reasoning dataset) that remain 95% unsolved indicates substantial headroom for future model development in complex reasoning domains.
C.2 How much has intelligence efficiency changed over time?
Figure 10: Perplexity and Accuracy per Joule Trends across Wildchat and NaturalReasoning.
Figure 10 presents results across single-turn chat and reasoning queries, evaluating intelligence efficiency across model-hardware pairs from April 2024 through August 2025. We measure both perplexity (left panel) and accuracy (right panel) normalized by energy consumption in joules per query, tracking nine distinct model families (Llama, Phi, Gemma, Mistral, Falcon, DeepSeek, Qwen, and GPT-OSS) deployed on various GPU configurations including NVIDIA A100 (40GB/80GB PCIe/SXM), H100 (80GB SXM), H200 (141GB HBM3e), and L40S (48GB) accelerators. Energy measurements capture end-to-end inference costs. The temporal snapshots at April 2024, August 2024, and August 2025 enable direct comparison of efficiency trajectories, revealing how successive generations of models and hardware migrate from suboptimal regions (high energy, low performance) toward optimal regions (low energy, high performance) as indicated by the shaded quadrants in each panel.
C.3 What efficiency gains can effective query routing deliver?
We compute cost per query using pricing available on OpenRouterAI (OpenRouter, 2025). Table 12 lists the token pricing used.
Model Input Cost (USD / 1M tokens) Output Cost (USD / 1M tokens)
Qwen3-4B 0.000 0.000
Qwen3-8B 0.035 0.138
Qwen3-14B 0.060 0.124
Qwen3-32B 0.100 0.450
Qwen3-235B 0.220 0.880
GPT-OSS-20B 0.03 0.14
GPT-OSS-120B 0.15 0.60
Table 12: Model pricing from OpenRouterAI. Costs are in USD per 1M tokens for input and output, as of August 2025.
C.4 How does model precision affect performance and efficiency?
Model quantization—reducing numerical precision from Fp16 to Fp8 or Fp4—decreases memory requirements and energy consumption during inference while introducing approximation error that may degrade model accuracy. To quantify this tradeoff, we evaluate eight open-source models from the Qwen3 and Gemma families across three precision levels: Fp16 (full precision), Fp8 (8-bit floating point), and Fp4 (4-bit floating point). For each model-precision pair, we measure accuracy on three reasoning-focused datasets: NaturalReasoning ($N=10,000$), SuperGPQA ($N=10,000$), and MMLU Pro ($N=10,000$).
Figure 11 shows that quantization from Fp16 to Fp4 yields energy reductions of $3-3.5\times$ with accuracy degradation of approximately 2.5 percentage points per precision step across all models and datasets. For example, on SuperGPQA, Qwen3-14B achieves 54.5% (Fp16), 52.0% (Fp8), and 49.0% (Fp4)—a total degradation of 5.5 percentage points despite a $3.23\times$ reduction in energy consumption. Larger models maintain their relative performance advantage even at lower precision: Qwen3-14B at Fp4 (49.0% accuracy) outperforms Qwen3-4B at Fp16 (48.5% accuracy) on SuperGPQA, indicating that model scale matters more than precision for reasoning tasks. These results demonstrate that Fp8 and Fp4 quantization enable practical deployment of local models with predictable performance tradeoffs, allowing system designers to select precision levels based on application-specific requirements while capturing most of the energy savings identified in Section 5.3.
Figure 11: Minimal Accuracy Degradation Shifting from FP16 to FP4 for Open-Source Local Models: Evaluation across three reasoning datasets ($N=10,000$ each) shows $2-3\%$ accuracy loss per precision step, demonstrating that FP8FP8/$FP4$ quantization enables efficient deployment with acceptable performance tradeoffs.
C.5 How do SOTA open-source LMs compare to the SOTA closed-source LMs on Chat and Reasoning Queries?
To evaluate the competitiveness of open-source models for local deployment, we compare the performance of state-of-the-art open-source models against leading closed-source models across single-turn chat and reasoning queries. We evaluate three closed-source frontier models—GPT-5-2025-08-07, Gemini-2.5-Pro, and Claude-Sonnet-4-5—against eight open-source models ranging from Qwen3-8B to Qwen3-235B-A22B, measuring performance on NaturalReasoning, MMLU Pro, and SuperGPQA. Table 13 shows that the best open-source model (Qwen3-235B-A22B) achieves 71.8% average accuracy across benchmarks, trailing the best closed-source model (GPT-5-2025-08-07, 77.9%) by 6.1 percentage points.
Performance gaps vary substantially by task type, as shown in Table 14. On MMLU Pro and SuperGPQA, open-source models nearly match closed-source performance: Qwen3-235B-A22B achieves 82.3% versus 87.4% on MMLU Pro (5.1% gap) and 63.1% versus 66.5% on SuperGPQA (3.4% gap). However, on NaturalReasoning, the gap widens to 12.9% (70.0% versus 82.9%), indicating that closed-source models maintain a substantial advantage on naturalistic reasoning tasks. Local models with deployment constraints ($\leq 20B$ active parameters) face larger gaps: the best local model (Qwen3-14B) trails closed-source models by 11.8–13.2% across benchmarks, with the smallest gap on MMLU Pro (11.8%) and the largest on NaturalReasoning (13.2%). These results demonstrate that while open-source models at frontier scale (235B parameters) approach closed-source performance on knowledge and reasoning benchmarks, practical local deployment using smaller models (14B parameters) requires accepting 11–13% accuracy degradation relative to closed-source alternatives.
Model Type WildChat NaturalReasoning MMLU Pro SuperGPQA Average
gpt-5-2025-08-07 Closed-Source 81.9% 82.9% 86.5% 64.4% 78.9%
gemini-2.5-pro Closed-Source 89.5% 77.9% 87.4% 66.5% 80.3%
claude-sonnet-4-5 Closed-Source 88.1% 76.9% 86.4% 60.1% 77.9%
Qwen/Qwen3-235B-A22B (Best OSS) Open-Source N/A 70.0% 82.3% 63.1% 71.8%
Qwen/Qwen3-32B Open-Source 76.1% 69.7% 77.9% 56.5% 70.1%
openai/gpt-oss-120b Open-Source 89.2% 65.0% 78.3% 55.3% 72.0%
Qwen/Qwen3-30B-A3B Open-Source 47.3% 64.3% 76.9% 57.4% 61.5%
Qwen/Qwen3-14B Open-Source 48.9% 60.0% 75.6% 56.2% 60.2%
openai/gpt-oss-20b Open-Source 77.3% 67.3% 73.4% 48.9% 66.7%
Qwen/Qwen3-8B Open-Source 50.2% 57.9% 73.3% 51.8% 58.3%
Table 13: Model performance comparison across benchmarks. Scores represent performance on Wildchat, NaturalReasoning, MMLU Pro, and SuperGPQA benchmarks. Qwen3-235B-A22B is used as the reference model for Wildchat evaluation.
Metric WildChat NaturalReasoning MMLU Pro SuperGPQA
Closed Best 89.5% 82.9% 87.4% 66.5%
Best Closed Model gemini-2.5-pro gpt-5-2025-08-07 gemini-2.5-pro gemini-2.5-pro
Open Best 89.2%* 70.0% 82.3% 63.1%
Best Open Model openai/gpt-oss-120b Qwen/Qwen3-235B-A22B Qwen/Qwen3-235B-A22B Qwen/Qwen3-235B-A22B
Gap $0.3\%$ $12.9\%$ $5.1\%$ $3.4\%$
Local Best ($\leq 20B$ Active) 89.2% 67.3% 80.3% 50.5%
Best Local Model openai/gpt-oss-120b openai/gpt-oss-20b openai/gpt-oss-120b Qwen/Qwen3-14B
Local Gap $0.3\%$ $5.6\%$ $7.1\%$ $16.0\%$
Table 14: Closed-Source vs. Open-Source Performance Gap by Task. Comparison between closed-source, open-source (all sizes), and local ($\leq 20B$ active parameters) models. *Excludes Qwen3-235B-A22B (reference model for WildChat).
C.6 How does performance on chat and reasoning queries connect to U.S. GDP?
To assess the economic relevance of local model performance improvements, we compute GDP-weighted accuracy for each model by weighting its performance on each economic category by that sector’s contribution to the 2024 U.S. GDP of \$29.18 trillion (U.S. Bureau of Economic Analysis, 2024). This metric attempts to quantify what proportion of economic value is relevant and addressable by local LMs, given the local model performances across single-turn chat and reasoning queries. Figures 12 and 13 shows that model improvements translate directly into expanded GDP coverage: on SuperGPQA, Qwen3-235B-A22B achieves 59.2% accuracy covering \$9.3T in relevant GDP (31.9% of total U.S. GDP), while on MMLU Pro, it reaches 84.5% accuracy covering \$7.6T (26.0% of total U.S. GDP). The strong positive correlation between model accuracy and GDP coverage across both benchmarks demonstrates that scaling model capabilities systematically expands the set of economically valuable tasks that can be automated.
Task type substantially affects GDP coverage: chat tasks in Wildchat show the highest coverage with gpt-oss-120b reaching 89.2% accuracy and covering \$20.3T (69.6% of U.S. GDP), while reasoning tasks in NaturalReasoning show lower coverage with Qwen3-235B-A22 achieving 69.3% accuracy but only covering \$6.8T (23.3% of U.S. GDP). This disparity reveals that current models excel at creative and conversational tasks that dominate economic activity, but struggle with technical reasoning tasks concentrated in specialized sectors like architecture, engineering, and physical sciences. The gap between chat coverage (69.6%) and reasoning coverage (23.3%) represents both a limitation of open-source local models and an economic opportunity: improving reasoning capabilities could unlock an additional \$13.5T in GDP-relevant tasks, suggesting that advances in technical reasoning would have substantial economic impact beyond current model capabilities.
Figure 12: Open-Source Local LMs Performance vs. U.S. GDP - WildChat and Natural Reasoning: Model accuracy on WildChat and Natural Reasoning benchmarks plotted against relevant GDP in trillions of dollars. Both benchmarks show continued performance improvements as training compute scales across models from Qwen3B-4B to Qwen3B-A22B-235B. For our calculations, we compute the weighted sum of an LM’s accuracy on each U.S. Labor category vs. the U.S. GDP associated with that category.
Figure 13: Open-Source Local LMs Performance vs. U.S. GDP - SuperGPQA and MMLU Pro: Model accuracy on SuperGPQA and MMLU Pro benchmarks plotted against relevant GDP in trillions of dollars. Both benchmarks show continued performance improvements as training compute scales across models from Qwen3B-4B to Qwen3B-A22B-235B. For our calculations, we compute the weighted sum of an LM’s accuracy on each U.S. Labor category vs. the U.S. GDP associated with that category.